# Scalable Human-Machine Point Cloud Compression

### 19 Feb 2024 v1

@@ II Related work
Specialized compression for machine tasks — often referred to as coding for machines — has been explored for images [17] and video [18] and, more recently, for point clouds [3]. Moreover, scalable multi-task coding approaches [19, 20] have shown that one can perform a machine vision task at a fairly low bitrate, while enabling other tasks, such as input reconstruction for human viewing, with an additional enhancement bitstream.
@@ III-C Proposed architecture
Our complete proposed codec architecture, based on PointNet++ [5], is shown in fig. 2, and the details of each block are shown in fig. 3. The input point cloud of $P$ points is represented as a matrix ${\boldsymbol{x}}\in\mathbb{R}^{3\times P}$ of $(x,y,z)$ coordinates. ${\boldsymbol{u}}$ is an additional set of attribute features (e.g., normals, color, etc.) that may be optionally compressed as well. At the beginning of the encoder, ${\boldsymbol{x}}$ and ${\boldsymbol{u}}$ are concatenated along the channel dimension so that ${\boldsymbol{u}}^{(0)}=\operatorname{concat}[{\boldsymbol{x}},{\boldsymbol{u}}]$. This ensures that the same encoding capabilities are available for both. We feed ${\boldsymbol{x}}^{(0)}={\boldsymbol{x}}$ and ${\boldsymbol{u}}^{(0)}$ into a sequence of downsampling blocks, which recursively downsample the point cloud into a smaller and smaller set of centroids.
@@
Each downsampling block is a PointNet++ “set abstraction” layer (see [5] for more details), with minor modifications. We used single-scale grouping (SSG) of points for our proposed codec, though it may likely be improved with multi-scale grouping (MSG) or multi-resolution grouping (MRG) as described in [5]. The “set abstraction” layer takes as input a (subsampled) point cloud ${\boldsymbol{x}}^{(i-1)}$ of shape $3\times P^{(i-1)}$ and a feature vector ${\boldsymbol{u}}^{(i-1)}$ of shape $D^{(i-1)}\times P^{(i-1)}$ containing information about each corresponding point. Using farthest point sampling (FPS), a set of $P^{(i)}$ centroids ${\boldsymbol{x}}^{(i)}$ of shape $3\times P^{(i)}$ is selected from ${\boldsymbol{x}}^{(i-1)}$. Then, the points in ${\boldsymbol{x}}^{(i-1)}$ are grouped into $P^{(i)}$ groups of $S^{(i)}$ points each. Then for each centroid, a ball query [^2] Note that due to the ball query, the same point may be assigned to multiple centroids. Also, centroids are not always present within their own group of points. Nonetheless, the ball query has the benefit of scale-invariant grouping. [5] is performed to find the first $S^{(i)}$ points that are within a radius of $R^{(i)}$ from the given centroid. The relative positions of each group of points is then computed with respect to its associated centroid, resulting in the residuals $r^{(i)}=\operatorname{ball\_query}({\boldsymbol{x}}^{(i-1)},{\boldsymbol{x}}^{(i)},R^{(i)})-\operatorname{repeat}({\boldsymbol{x}}^{(i)},S^{(i)})$ of shape $3\times P^{(i)}\times S^{(i)}$. Each point in ${\boldsymbol{r}}^{(i)}$ is concatenated with its respective feature vector in ${\boldsymbol{u}}^{(i-1)}$, resulting in a grouped feature tensor ${\boldsymbol{u_{g}}}^{(i-1)}$ of shape $(D^{(i-1)}+3)\times P^{(i)}\times S^{(i)}$. Then, ${\boldsymbol{u_{g}}}^{(i-1)}$ is fed into a miniature group-wise “PointNet encoder” block consisting of a shared multi-layer perceptron (MLP) with a max pooling operation at the end. That is, the same PointNet encoder is applied to each group of features independently and identically. This finally results in a feature vector ${\boldsymbol{u}}^{(i)}$ of shape $D^{(i)}\times P^{(i)}$, from which the previous centroid point cloud ${\boldsymbol{x}}^{(i-1)}$ may be reconstructed.
Three downsampling blocks are used in our proposed codec. The last downsampling block groups all the remaining points into a single group, and thus $P^{(3)}=1$.
@@
At each level $i<3$, the grouped features ${\boldsymbol{u_{g}}}^{(i)}$ are fed into a standard transform-encoder-decoder-transform compression pipeline. That is, the latent representation ${\boldsymbol{y}}^{(i)}={h_{a}}^{(i)}({\boldsymbol{u_{g}}}^{(i)})$ of shape $M^{(i)}\times P^{(i)}$ is computed, then quantized as ${\boldsymbol{\hat{y}}}^{(i)}=Q^{(i)}({\boldsymbol{y}}^{(i)})$, and then entropy-coded using a learned entropy model. The resulting bitstream is transmitted and then decoded as ${\boldsymbol{\hat{y}}}^{(i)}$. Then, the decoder-side transform $h_{s}^{\,(i)}$ is applied to obtain a set of grouped features ${\boldsymbol{\hat{u}_{g}}}^{(i)}=h_{s}^{\,(i)}({\boldsymbol{\hat{y}}}^{(i)})$ used for upsampling. (Note that ${\boldsymbol{\hat{u}_{g}}}^{(i)}$ and ${\boldsymbol{u_{g}}}^{(i)}$ may inhabit entirely different feature spaces, and are unrelated.)
@@
After the last downsampling block, the feature vector ${\boldsymbol{u}}^{(3)}$ is fed into the scalable multi-task compression pipeline described in section III-B. The latent representation ${\boldsymbol{y}}^{(3)}={h_{a}}^{(3)}({\boldsymbol{u}}^{(3)})$ is computed, then quantized as ${\boldsymbol{\hat{y}}}^{(3)}=Q^{(3)}({\boldsymbol{y}}^{(3)})$, and then entropy-coded to generate a base bitstream for ${\boldsymbol{\hat{y}}}_{1}^{(3)}$ and an enhancement bitstream for ${\boldsymbol{\hat{y}}}_{2}^{(3)}$. The base bitstream is decoded as ${\boldsymbol{\hat{y}}}_{1}^{(3)}$, and then fed into the decoder-side transform $h_{s}^{(3)}$ to obtain the class prediction ${\boldsymbol{\hat{t}}}=h_{s}^{(3)}({\boldsymbol{\hat{y}}}_{1}^{(3)})$. The enhancement bitstream is decoded as ${\boldsymbol{\hat{y}}}_{2}^{(3)}$, and then its (detached) concatenation with ${\boldsymbol{\hat{y}}}_{1}^{(3)}$ is fed into the decoder-side transform $h_{s}^{(3)}$ to obtain ${\boldsymbol{\hat{u}_{g}}}^{(3)}=h_{s}^{(3)}(\operatorname{concat}[\operatorname{detach}({\boldsymbol{\hat{y}}}_{1}^{(3)}),{\boldsymbol{\hat{y}}}_{2}^{(3)}])$. Note that we split only ${\boldsymbol{y}}^{(3)}$ into ${\boldsymbol{y}}^{(3)}_{1}$ and ${\boldsymbol{y}}^{(3)}_{2}$. This is because ${\boldsymbol{y}}^{(3)}_{1}$ suffices for the classification task, while ${\boldsymbol{y}}^{(3)}_{2}$ and ${\boldsymbol{y}}^{(j)}$ for $j<3$ are only needed for input reconstruction. Due to space constraints, in the experiments we only test input reconstruction with all enhancement layers, but the proposed architecture also allows for quality- or density-scalable input reconstruction.
@@
To perform input reconstruction, all enhancement bitstreams are decoded into their respective latent representations, from which various ${\boldsymbol{\hat{u}_{g}}}^{(i)}$ are obtained. These are upscaled via a sequence of upsampling blocks, while concatenating the features at each level appropriately. An upsampling block takes in $E^{(i+1)}+D^{(i)}+3\cdot(1-\delta_{i,3})$ channels and outputs $E^{(i)}$ channels, where $\delta_{i,j}$ is the Kronecker delta ($\delta_{i,j}=1$ if $i=j$, 00 otherwise). Each upsampling block consists of a simple two-layer point-wise MLP. For upsampling blocks at levels deeper than $i>0$, batch normalizations are also present. Furthermore, each such block is followed by a reshape-transpose-reshape combination that converts the resulting shape $(E^{(i)}\cdot S^{(i)})\times P^{(i)}$ into the output shape $E^{(i)}\times(S^{(i)}\cdot P^{(i)})$. To ensure compatibility of shapes among various components during reconstruction, we enforce the constraint $P^{(i-1)}=P^{(i)}\cdot S^{(i)}$.
Our codec is provided in “full” and “lite” configurations, as detailed in table I.
@@ IV Experiments
(a) Rate-accuracy for base task.
(b) Rate-accuracy for base task (zoomed in).
(c) Rate-distortion for reconstruction task.
Figure 4: Rate-accuracy and rate-distortion curves for the proposed codec on the ModelNet40 dataset (1 kibibit = 1024 bits).
@@ V Results
fig. 3(c) shows the rate-distortion curves for the reconstruction task in terms of the rate versus the Chamfer distance. For the proposed codec, the rate includes both the base and enhancement bitstreams. The figure shows that our proposed codec is also capable of achieving better reconstruction quality than conventional codecs at low rates, below 1.6 Kib. As the rate increases, the conventional codecs eventually catch up in terms of reconstruction quality.

### 22 Feb 2024 v2

@@ II Related work
Specialized compression for machine tasks — often referred to as coding for machines — has been explored for images [17] and video [18] and, recently, for point clouds [3]. Moreover, scalable multi-task coding approaches [19, 20] have shown that one can perform a machine vision task at a fairly low bitrate, while enabling other tasks, such as input reconstruction for human viewing, with an additional enhancement bitstream.
@@ III-C Proposed architecture
Our complete proposed codec architecture, based on PointNet++ [5], is shown in fig. 2, and the details of each block are shown in fig. 3. Our codec is provided in “full” and “lite” configurations, as detailed in table I. The input point cloud of $P$ points is represented as a matrix ${\boldsymbol{x}}\in\mathbb{R}^{3\times P}$ of $(x,y,z)$ coordinates. ${\boldsymbol{u}}$ is an additional set of attribute features (e.g., normals, color, etc.) that optionally may also be compressed. At the beginning of the encoder, ${\boldsymbol{x}}$ and ${\boldsymbol{u}}$ are concatenated along the channel dimension so that ${\boldsymbol{u}}^{(0)}=\operatorname{concat}[{\boldsymbol{x}},{\boldsymbol{u}}]$. This ensures that the same encoding capabilities are available for both. We feed ${\boldsymbol{x}}^{(0)}={\boldsymbol{x}}$ and ${\boldsymbol{u}}^{(0)}$ into a sequence of downsampling blocks. Each $i$-th downsampling block takes in ${\boldsymbol{x}}^{(i-1)}$ and ${\boldsymbol{u}}^{(i-1)}$, and outputs and a smaller set of centroids ${\boldsymbol{x}}^{(i)}$ and features ${\boldsymbol{u}}^{(i)}$, along with features ${\boldsymbol{u_{g}}}^{(i-1)}$ that are grouped by the centroids. The final ${\boldsymbol{u}}^{(3)}$ is compressed in a multi-task scalable manner, from which ${\boldsymbol{\hat{u}}}^{(3)}$ is derived. Each ${\boldsymbol{u_{g}}}^{(i-1)}$ is compressed using a standard transform-encoder-decoder-transform compression pipeline to obtain ${\boldsymbol{\hat{u}_{g}}}^{(i-1)}$. A sequence of upsampling blocks is then applied, where each $i$-th upsampling block takes in ${\boldsymbol{\hat{u}}}^{(i)}$. For $i>0$, the output of the $i$-th upsampling block is concatenated with ${\boldsymbol{\hat{u}_{g}}}^{(i-1)}$ to give ${\boldsymbol{\hat{u}}}^{(i-1)}$. For $i=0$, the output is ${\boldsymbol{\hat{x}}}$.
III-C1 Downsampling
@@
Each downsampling block is a PointNet++ “set abstraction” layer (see [5] for more details), with minor modifications. We used single-scale grouping (SSG) of points for our proposed codec, though it may likely be improved with multi-scale grouping (MSG) or multi-resolution grouping (MRG) as described in [5]. The “set abstraction” layer takes as input a (subsampled) point cloud ${\boldsymbol{x}}^{(i-1)}$ of shape $3\times P^{(i-1)}$ and features ${\boldsymbol{u}}^{(i-1)}$ of shape $D^{(i-1)}\times P^{(i-1)}$ containing information about each corresponding point. Using farthest point sampling (FPS), a set of $P^{(i)}$ centroids ${\boldsymbol{x}}^{(i)}$ of shape $3\times P^{(i)}$ is selected from ${\boldsymbol{x}}^{(i-1)}$. Then, the points in ${\boldsymbol{x}}^{(i-1)}$ are grouped into $P^{(i)}$ groups of $S^{(i)}$ points each. Then for each centroid, a ball query [^2] Note that due to the ball query, the same point may be assigned to multiple centroids. Also, centroids are not always present within their own group of points. Nonetheless, the ball query has the benefit of scale-invariant grouping. [5] is performed to find the first $S^{(i)}$ points that are within a radius of $R^{(i)}$ from the given centroid. The relative positions of each group of points is then computed with respect to its associated centroid, resulting in the residuals $r^{(i)}=\operatorname{ball\_query}({\boldsymbol{x}}^{(i-1)},{\boldsymbol{x}}^{(i)},R^{(i)})-\operatorname{repeat}({\boldsymbol{x}}^{(i)},S^{(i)})$ of shape $3\times P^{(i)}\times S^{(i)}$. Each point in ${\boldsymbol{r}}^{(i)}$ is concatenated with its respective feature vector in ${\boldsymbol{u}}^{(i-1)}$, resulting in a grouped feature tensor ${\boldsymbol{u_{g}}}^{(i-1)}$ of shape $(D^{(i-1)}+3)\times P^{(i)}\times S^{(i)}$. Then, ${\boldsymbol{u_{g}}}^{(i-1)}$ is fed into a miniature group-wise “PointNet encoder” block consisting of a shared multi-layer perceptron (MLP) with a max pooling operation at the end. That is, the same PointNet encoder is applied to each group of features independently and identically. This results in ${\boldsymbol{u}}^{(i)}$ of shape $D^{(i)}\times P^{(i)}$. Three downsampling blocks are used in our proposed codec. The last downsampling block groups all the remaining points into a single group, i.e., $P^{(3)}=1$.
III-C2 Feature compression
@@
At each level $i<3$, a latent representation ${\boldsymbol{y}}^{(i)}={h_{a}}^{(i)}({\boldsymbol{u_{g}}}^{(i)})$ of shape $M^{(i)}\times P^{(i)}$ is computed, then quantized as ${\boldsymbol{\hat{y}}}^{(i)}=Q({\boldsymbol{y}}^{(i)})$, and then entropy-coded using a learned entropy model. The resulting bitstream is transmitted and then decoded as ${\boldsymbol{\hat{y}}}^{(i)}$. From this, a set of grouped features ${\boldsymbol{\hat{u}_{g}}}^{(i)}=h_{s}^{\,(i)}({\boldsymbol{\hat{y}}}^{(i)})$ is computed for usage during upsampling. (Note that ${\boldsymbol{\hat{u}_{g}}}^{(i)}$ and ${\boldsymbol{u_{g}}}^{(i)}$ may inhabit entirely different feature spaces, and are unrelated.)
@@
The final feature vector ${\boldsymbol{u}}^{(3)}$ is fed into the scalable multi-task compression pipeline described in section III-B. The latent representation ${\boldsymbol{y}}^{(3)}={h_{a}}^{(3)}({\boldsymbol{u}}^{(3)})$ is computed, then quantized as ${\boldsymbol{\hat{y}}}^{(3)}=Q({\boldsymbol{y}}^{(3)})$, and then entropy-coded to generate a base bitstream for ${\boldsymbol{\hat{y}}}_{1}^{(3)}$ and an enhancement bitstream for ${\boldsymbol{\hat{y}}}_{2}^{(3)}$. The base bitstream is decoded as ${\boldsymbol{\hat{y}}}_{1}^{(3)}$, and then fed into a decoder-side task backend to obtain the class prediction ${\boldsymbol{\hat{t}}}=g_{s,t}^{(3)}({\boldsymbol{\hat{y}}}_{1}^{(3)})$. The enhancement bitstream is decoded as ${\boldsymbol{\hat{y}}}_{2}^{(3)}$, and then its concatenation with ${\boldsymbol{\hat{y}}}_{1}^{(3)}$ (detached) is fed into a decoder-side transform to obtain ${\boldsymbol{\hat{u}_{g}}}^{(3)}=h_{s}^{(3)}(\operatorname{concat}[\operatorname{detach}({\boldsymbol{\hat{y}}}_{1}^{(3)}),{\boldsymbol{\hat{y}}}_{2}^{(3)}])$.
Note that we split only ${\boldsymbol{y}}^{(3)}$ into ${\boldsymbol{y}}^{(3)}_{1}$ and ${\boldsymbol{y}}^{(3)}_{2}$. This is because ${\boldsymbol{y}}^{(3)}_{1}$ suffices for the classification task, while ${\boldsymbol{y}}^{(3)}_{2}$ and ${\boldsymbol{y}}^{(j)}$ for $j<3$ are only needed for input reconstruction. For brevity, in the experiments we only test input reconstruction with all enhancement layers, but the proposed architecture also allows for quality- or density-scalable input reconstruction.
III-C3 Upsampling
@@
An upsampling block takes in $E^{(i+1)}+D^{(i)}+3\cdot(1-\delta_{i,3})$ channels and outputs $E^{(i)}$ channels, where $\delta_{i,j}$ is the Kronecker delta ($\delta_{i,j}=1$ if $i=j$, 00 otherwise). Each upsampling block consists of a simple two-layer point-wise MLP. For upsampling blocks at levels deeper than $i>0$, batch normalizations are also present. Furthermore, each such block is followed by a reshape-transpose-reshape combination that converts the resulting shape $(E^{(i)}\cdot S^{(i)})\times P^{(i)}$ into the output shape $E^{(i)}\times(S^{(i)}\cdot P^{(i)})$. To ensure compatibility of shapes among various components during reconstruction, we enforce the constraint $P^{(i-1)}=P^{(i)}\cdot S^{(i)}$.
@@ IV Experiments
(a) RA for base task.
(b) RA for base task (zoomed in).
(c) RD for reconstruction task.
(d) RD for reconstruction task.
Figure 4: Rate-accuracy (RA) and rate-distortion (RD) curves on the ModelNet40 dataset, with rate units of bits per point (bpp) scaled for 1024 points.
@@ V Results
figs. 3(c) and 3(d) show the rate-distortion curves for the reconstruction task in terms of the rate versus the Chamfer distance and D1-PSNR [27]. For the proposed codec, the rate includes both the base and enhancement bitstreams. The figures show that our proposed codec is also capable of achieving better reconstruction quality than conventional codecs at low rates, below 1.6 bpp. As the rate increases, the conventional codecs eventually catch up in terms of reconstruction quality.